OpenStudy (anonymous):

The Radius of a Copper atom is 128 pm. What is the diameter (in micrometer), circumference (m), and area (km^2) of it?

OpenStudy (cuanchi):

you need the following formulas and conversion factors, where r= radius; π=3.14) \[diameter=2 r\] \[circunference=2\pi r\] \[area=4\pi r^2\] \[\frac{ 1\mu m }{10^6 p m }\] \[\frac{ 1\ m }{10^{12} p m }\] \[\left( \frac{ 1km }{10^3m } \times \frac{ 1m }{10^{12} p m }\right)^2=\frac{ 1km }{ 10^{30}p m }\]

OpenStudy (anonymous):

a)\[2.56 x 10^{-7}\] b) \[2.56 x 10^{-10}\] c) \[5.12 x 10^{-13}\] Is that right, @Cuanchi ?

OpenStudy (cuanchi):

NOOO!!! a) 2.56x10−4 b) 8.04x10-10 c) 2.06x10-10